OpenStudy (anonymous):

A dispenser holds 5652 cm3 of liquid soap and is now full. The radius of the dispenser is 7.5 cm. What is the difference between the height of the soap in the full dispenser and the height when 4239 cm3 of soap remains in the dispenser? Use 3.14 to approximate pi.

OpenStudy (anonymous):

POP THE MOLLY.....IM.....SWEATIN.......THANK YOU

OpenStudy (anonymous):

BYE

OpenStudy (anonymous):

@Data_LG2

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

First, what is the shape of the dispenser?

OpenStudy (anonymous):

let's assume it's a cylinder. Now, what's the formula to get the volume of a cylinder?

OpenStudy (anonymous):

\(\sf V_{cylinder}= \pi r^2h\) now we are looking for the height. let's rearrange this equation in terms of h. dividing both sides by pi r squared... you'll get \(\sf \Large h=\frac{ V_{cylinder}}{\pi r^2}\)

OpenStudy (anonymous):

Next, solve the value of h when ~the dispenser is full, \(\sf V_{cylinder}=5652\ cm^3,\ r= 7.5,\ \pi=3.14\) and when ~the dispenser is not full \(\sf V_{cylinder}=4239\ cm^3,\ r= 7.5,\ \pi=3.14\)

OpenStudy (anonymous):

Lastly, after you find those values of h, subtract them.

OpenStudy (anonymous):

follow these steps and tell me what you get for each step and I'll check it later. any confusions, don't hesitate to ask :)

OpenStudy (anonymous):

for when the dispenser is full i got 90714.69 but im not sure its right and when it isnt full i got 99828.45 but the numbers didnt seem right to me but idk

OpenStudy (anonymous):

@Data_LG2

OpenStudy (anonymous):

okay, let see  is your solution the same ?